38. An increase in the supply of labor с.increases the value of ~ginal Product of er and enes D. decreases the value of the marginal product of labor and increases the wage the wage. 39. A decrease in the demand for fish A. decreases the value of the marginal product of fishemen reduces their wage, and reduces employment in the fishing industry employment in the fishing industry employment in the fishing industry employment in the fishing industry B. increases the value of the marginal product of fishermen, increases their wage, and increases C. decreases the value of the marginal product of fishermen, reduces their wage, and increases D. increases the value of the marginal product of fishermen, increases their wage, and decreases 40. In a competitive labor market, the number of workers gets hired where wage equals marginal revenue A. True B. False For questions 41, 42, and 43, use the table shown bellow Labor (number 2 3 12 14 15 5 41. What is the marginal product of labor for the fourth worker? A. 0 B. 2 C. 12 D. 14 42. If the price of output is $4 per unit, what is the marginal revenue product for the fourth worker in a competitive labor market? A. $4 B. $8 C. $12 D. $56 43. If this profit-maximizing firm sells its output in a competitive market for $4 per unit and hires labor in a competitive market for $8 per hour, then this firm should hire A. one worker. B. two workers C. three workers. D. four workers. 44. The opportunity cost of leisure: A. increases as wages get higher B. decreases as wages get higher C. remains the same as wages get higher. D. has nothing to do with wages
Answer #1

38. B
(As supply of labor increases, marginal product of labor declines. Thus, both value of marginal product of labor and wage rate declines)

39. A
(As demand fo fish decline so does the demand for fishermen will decline. Thus, value of marginal product of labor and wage rate decline with less employment)

40. True
(A competitive firm hires workers upto the point where wage = marginal revenue product so that profits are maximized)

41. B. 2
(Marginal product of fourth worker = (output of 4th worker-output of 3rd worker)/Change in workers = (14-12)/(4-3) = 2/1 = 1)

Note: Post four MCQs at a time.